# A Survey on LLM-as-a-Judge

### v4

@@ 4.3. Adversarial Robustness
Research (Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Similarly, work by Lee et al. (Lee et al., 2024a) introduced EMBER, a benchmark that revealed biases in when assessing outputs with epistemic markers, such as expressions of certainty or uncertainty. Furthermore, other work (Zheng et al., 2024a) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. Several recent works (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90% believe this is better”. Other research (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
5. Effectiveness Experiments
@@
In Section 3, we have introduced the improvement strategies adopted by researchers in existing LLM-as-a-judge works to improve the evaluation capabilities of LLM. Although numerous works have proposed meta-evaluation benchmarks to assess the performance of LLMs in evaluation tasks, as shown in Table 1, there is still a lack of meta-evaluation of wh ether these improvement strategies effectively optimize LLM evaluators and which dimensions of evaluation performance are being enhanced. Some improvement strategies may fail to improve the LLM evaluators’ performance or mitigate biases in practical use, leading to computing waste. In this section, based on the benchmarks mentioned in Section 4, we designed a robust and scalable meta-evaluation tool as shown in Fig. 13, and conducted a simple meta-evaluation experiment on the improvement strategies summarized in Section 3, examining their effectiveness from the perspectives of biases and agreement with human evaluation.
@@ 5.1.1. Evaluation Dimensions and Benchmarks
The most direct metric to reflect the quality of automatic evaluation is the alignment with human evaluation. We use LLMEval2 (Zhang et al., 2023c) to assess the alignment of LLM-as-a-judge with human evaluations. LLMEval2 is the largest and most diverse evaluation benchmark for LLM-as-a-judge to date, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Each sample consists of a question, a pair of candidate responses, and a human label indicating the preferred response.
@@ 5.2.1. Comparison with Different LLMs
LLMs Alignment Biases with Position Length Concre- Empty Content Nested Familiar Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-4-turbo 61.54 80.31 91.18 89.29 65.38 95.83 70.83 100.0 GPT-3.5-turbo 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 Qwen2.5-7B-Instruct 56.54 63.50 64.71 71.43 69.23 91.67 45.83 83.33 LLaMA3-8B-Instruct 50.72 38.85 20.59 57.14 65.38 75.00 45.83 54.17 Mistral-7B-Instruct-v0.3 55.42 59.78 26.47 67.86 53.85 66.67 37.50 41.67 Mixtral-8$\times$7B-Instruct-v0.1 56.29 59.06 50.00 78.57 42.31 83.33 29.17 83.33 gemini-2.0-thinking 60.75 76.84 94.12 89.29 50.00 100.00 83.33 100.00 o1-mini 60.16 76.73 91.18 89.29 53.85 95.83 75.00 95.83 deepseek r1 56.48 69.17 94.12 100.00 50.00 100.00 75.00 87.50
@@
LLMs human=model1 human=model2 human=TIE (n=2071) (n=2071) (n=964) aligned accuracy (%) aligned accuracy (%) aligned accuracy (%) GPT-4-turbo 1418 68.47 1438 69.44 286 29.67 gemini-2.0-thinking 1354 65.38 1621 78.27 127 13.17 o1-mini 1444 69.72 1401 67.65 227 23.55 deepseek r1 1369 66.10 1342 64.80 173 17.95
Table 3. The evaluation results of each human label in LLMEval2.
@@ 5.2.3. Evaluation of Reasoning LLM-as-a-Judge
As discussed in Sections 2.4 and 2.5, judgment serves as the foundation for effective reasoning capabilities. In other words, models with stronger reasoning capabilities are generally better equipped to perform as reliable judges. To validate this assumption, we conducted evaluations on several reasoning LLMs including o1-mini, Gemini-thinking and Deepseek-R1. The results in Tables 2 and 3 provide key insights into the performance of reasoning-focused LLMs. While these models—gemini-2.0-thinking, o1-mini, and deepseek r1—demonstrate competitive alignment and accuracy relative to the top-performing GPT-4-turbo, their improvements in tasks requiring human alignment are not as pronounced as expected.
GPT-4-turbo remains the benchmark for alignment, achieving the highest accuracy rates of 68.47% and 69.44% in the human=model1 and human=model2 scenarios, respectively. It also excels in resolving ambiguous cases, with an accuracy of 29.67% in the human=TIE scenario, outperforming all other models. Among reasoning-enhanced models, gemini-2.0-thinking shows strong performance in the human=model2 scenario, achieving an accuracy of 78.27%—surpassing GPT-4-turbo—highlighting its capability to identify models that align more closely with human judgment. However, this strength does not consistently extend to other tasks or scenarios. Similarly, o1-mini and deepseek-R1, while trailing slightly behind GPT-4-turbo and gemini-2.0-thinking, outperform non-GPT models like Mixtral-8$\times$7B-Instruct-v0.1, demonstrating the added value of reasoning-based enhancements in alignment tasks. These results indicate that reasoning-enhanced LLMs provide meaningful advancements over baseline models but fall short of delivering consistent advantages in alignment-related tasks, suggesting room for further optimization in this area.
@@ 6. LLM-as-a-Judge and o1-like Reasoning Enhancement
When faced with a challenging problem, humans often spend considerable time and effort reflecting on various possibilities before arriving at a solution. In a similar fashion, o1, an advanced model developed by OpenAI, engages in a structured chain of thought to solve complex tasks. This process of deliberate reasoning allows o1 to continuously refine its approach (Qin et al., 2024c), step by step, as it navigates through difficult scenarios. A key factor in enhancing o1’s reasoning is the integration of LLM-as-a-Judge, which evaluates the model’s reasoning paths at each stage. As o1 works through a problem, the judge provides feedback that helps the model improve by pointing out inconsistencies, suggesting corrections, and identifying simpler ways to break down difficult tasks. By using the feedback from its own evaluations, much like a constitutional AI framework, o1 is able to adapt its reasoning strategies and enhance its performance. Through reinforcement learning, o1 fine-tunes its strategies, learning not only from its successes but also from its mistakes. The combination of LLM-as-a-Judge, reinforcement learning, and the feedback loop from constitutional evaluation allows o1 to dynamically adjust its reasoning, ensuring that the model continuously improves its ability to solve complex problems over time. This synergy between reasoning and judgment, combined with continuous feedback, drives o1’s advanced problem-solving capabilities.
@@
Reasoning, as a cognitive process, involves applying logic and evidence to derive conclusions. It is central to intellectual tasks such as decision-making, problem-solving, and critical analysis. Reasoning requires evaluating multiple possibilities and determining the most logically sound and coherent path. In contrast, LLM-as-a-Judge refers to the use of LLMs to perform judgment tasks, such as evaluating, scoring, ranking, or selecting optimal answers based on generated outputs. This concept parallels a judge’s role in ensuring fairness, accuracy, and coherence in competitive contexts. Although reasoning and judgment are separate concepts, they are closely connected. Reasoning frequently relies on judgment to evaluate intermediate steps, improve logic, and guarantee clarity in its results. When the process involves an infinite number of judgments, we can consider it as a process that approximates both reasoning and thinking. At the same time, effective judgment relies on strong reasoning capabilities to evaluate options against a set of logical criteria, which we discussed in Section 2.5 and Section 5.2.3. Therefore, LLM-as-a-Judge not only evaluates outputs but also enhances the reasoning process by helping to identify the most coherent, accurate solutions.
@@ 7. Applications
LLMs’ abilities as evaluators have gained widespread recognition in specialized fields, especially in complex, qualitative areas like legal texts, mathematical reasoning, and scientific research. This section reviews recent developments in LLM-as-a-judge applications across finance, law, science, and other industries, investigating how domain knowledge and LLM evaluators can further expand their impact in critical areas.
@@ 7.1.1. NLP
Text Generation Text generation tasks, such as dialog response generation, summarization, story creation, and creative writing, require content that is safe, accurate, and contextually relevant, though there isn’t a single ”correct” answer (Badshah and Sajjad, 2024; Bermejo, 2024). Unlike traditional metrics-based evaluations, LLM-as-a-judge offers a nuanced, adaptable, and customized assessment. According to Zheng et al. (2023d), LLMs like GPT-4 can evaluate text generation comparably to humans. This method has been used to evaluate outputs from single models and to compare multiple models in competitive settings. For instance, Gao et al. (2023c) employ ChatGPT for human-like summarization evaluation, while Wu et al. (2023) propose a comparison-based framework where LLMs act as judges to evaluate summarization quality.
@@
Reasoning Enhancing the reasoning capabilities of LLMs can overcome the limitations of scaling laws, unlocking their full potential. Effective reasoning is essential for tackling complex problems, making informed decisions, and delivering accurate, context-aware responses. Wei et al. (2022b) introduce Chain-of-Thought (CoT) prompting to facilitate step-by-step reasoning. More sophisticated cognitive structures (Yao et al., 2023a; Hao et al., 2023b) have been proposed to further enhance reasoning, yet selecting a reliable reasoning path remains a significant challenge. LLM-as-a-judge has been employed to address this issue.
@@
Some studies focus on sample-level reasoning path selection. Gao et al. (2023a) present a strategy evaluator for assessing candidate strategies. Kawabata and Sugawara (2024) propose REPS (Rationale Enhancement through Pairwise Selection), which uses pairwise self-evaluation to select valid rationales. Lahoti et al. (2023) demonstrate that LLMs can identify and enhance response diversity by aggregating multiple critiques. In multi-agent frameworks, Liang et al. (2023) introduce multi-agent debating (MAD), where a judge LLM selects the most reasonable response. Similarly, Li et al. (2024d) utilize a judge LLM in layer-based multi-agent collaboration to improve response quality and efficiency.
@@
For step-level reasoning path selection, LLMs act as process reward models (PRMs) to evaluate state scores. Creswell et al. (2023) break down reasoning into Selection and Inference, using LLMs to judge potential reasoning traces. Xie et al. (2024b) propose the Kwai-STaR framework, transforming LLMs into state-transition reasoners for mathematical reasoning. Lightman et al. (2023b) train LLMs as PRMs for inference-time supervision and best-of-N sampling. Setlur et al. (2024) introduce process advantage verifiers (PAVs) to generate rewards based on the likelihood of future correct responses. Advanced cognitive structures are also simulated; Hao et al. (2023b) use LLMs as a world model with Monte Carlo Tree Search (MCTS) for deliberate path selection. Besta et al. (2024) model LLM outputs as graphs to evaluate coherence and logical reasoning. Additionally, critique-based LLM judges (Ankner et al., 2024; Lan et al., [n. d.]; Yao et al., 2024; Yuksel and Sawaf, 2024) provide detailed feedback to enhance the reasoning process.
@@
Yao et al. (2023b) pioneered the use of LLMs in an interleaved manner to generate reasoning traces and task-specific actions. Reasoning traces guide the model in updating action plans, while actions facilitate interaction with external sources. Building on this, Yang et al. (2023) introduced Auto-GPT, which leverages LLM-as-a-judge to enhance tool usage accuracy. By integrating a variety of external tools, LLMs become more versatile, improving planning performance through judicious tool selection. Sha et al. (2023) explored the potential of LLMs in decision-making for complex autonomous driving scenarios, requiring human-like commonsense reasoning. Zhou et al. (2024) employed a self-discovery process where LLMs judge queries and select the most suitable reasoning structure for subsequent inference.
@@
Retrivel The role of LLM-as-a-judge in retrieval encompasses both traditional document ranking and dynamic Retrieval-Augmented Generation (RAG) approaches. In traditional retrieval, LLMs enhance ranking accuracy through advanced prompting techniques, enabling effective document ordering with minimal labeled data. RAG frameworks leverage LLMs’ ability to generate content guided by retrieved information, supporting applications requiring complex or evolving knowledge integration.
@@
Recent studies have explored LLMs as judges for document ranking, aiming to boost precision and reduce reliance on extensive training data. Zhuang et al. (2024a) embed fine-grained relevance labels within LLM prompts, enabling models to distinguish subtle relevance variations for refined document ordering. Innovations in listwise ranking include Ma et al. (2023)’s Listwise Reranker with a Large Language Model (LRL), which reorders document identifiers without task-specific training data. Zhuang et al. (2024b) introduce a Setwise prompting strategy for zero-shot ranking, enhancing efficiency without sacrificing performance. To address positional biases, Tang et al. (2024b) propose permutation self-consistency, averaging multiple list orders to yield order-independent rankings. Qin et al. (2024b) critique pointwise and listwise ranking prompts, proposing Pairwise Ranking Prompting (PRP) with medium-sized, open-source LLMs as a cost-efficient alternative to larger models.
@@
Domain-specific RAG systems reveal LLMs’ potential to navigate complex queries by integrating specialized knowledge structures. Wang et al. (2024d) present BIORAG, enhancing vector retrieval with hierarchical knowledge structures and a self-aware evaluated retriever. Li et al. (2024e) introduce DALK, combining an LLM with a continuously evolving Alzheimer’s Disease knowledge graph, using self-aware knowledge retrieval for noise filtering. Jeong et al. (2024) propose Self-BioRAG, adapting RAG principles to biomedical applications, with LLMs selecting the best evidence for answer generation.

### v5

@@ 4.3. Adversarial Robustness
Research (Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Similarly, work by Lee et al. (Lee et al., 2024a) introduced EMBER, a benchmark that revealed biases in when assessing outputs with epistemic markers, such as expressions of certainty or uncertainty. Furthermore, other work (Zheng et al., 2024a) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. Several recent works (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90 Other research (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
5. Meta-evaluation Experiment
@@
In Section 3, we have introduced the improvement strategies adopted by researchers in existing LLM-as-a-judge works to improve the evaluation capabilities of LLM. Although numerous works have proposed meta-evaluation benchmarks to assess the performance of LLMs in evaluation tasks, as shown in Table 1, there is still a lack of meta-evaluation of wh ether these improvement strategies effectively optimize LLM evaluators and which dimensions of evaluation performance are being enhanced. Some improvement strategies may fail to improve the LLM evaluators’ performance or mitigate biases in practical use, leading to computing waste.
In this section, based on the benchmarks mentioned in Section 4, we designed a robust and scalable meta-evaluation tool as shown in Fig. 14, and conducted a simple meta-evaluation experiment on the improvement strategies summarized in Section 3, examining their effectiveness from the perspectives of biases and agreement with human evaluation.
@@ 5.1.1. Evaluation Dimensions and Benchmarks
The most direct metric to reflect the quality of automatic evaluation is the alignment with human evaluation. We use LLMEval2 (Zhang et al., 2023d) to assess the alignment of LLM-as-a-judge with human evaluations. LLMEval2 is the largest and most diverse evaluation benchmark for LLM-as-a-judge to date, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Each sample consists of a question, a pair of candidate responses, and a human label indicating the preferred response.
@@ 5.2.1. Comparison with Different LLMs
LLMs Alignment Biases with Position Length Concre- Empty Content Nested Familiar Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-4-turbo 61.54 80.31 91.18 89.29 65.38 95.83 70.83 100.0 GPT-3.5-turbo 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 Qwen2.5-7B-Instruct 56.54 63.50 64.71 71.43 69.23 91.67 45.83 83.33 LLaMA3-8B-Instruct 50.72 38.85 20.59 57.14 65.38 75.00 45.83 54.17 Mistral-7B-Instruct-v0.3 55.42 59.78 26.47 67.86 53.85 66.67 37.50 41.67 Mixtral-8$\times$7B-Instruct-v0.1 56.29 59.06 50.00 78.57 42.31 83.33 29.17 83.33 gemini-2.0-thinking 60.75 76.84 94.12 89.29 50.00 100.00 83.33 100.00 o1-mini 60.16 76.73 91.18 89.29 53.85 95.83 75.00 95.83 o3-mini 61.66 74.63 82.35 92.86 73.08 95.83 87.50 91.67 deepseek r1 56.48 69.17 94.12 100.00 50.00 100.00 75.00 87.50
@@
LLMs human=model1 human=model2 human=TIE aligned / total accuracy (%) aligned / total accuracy (%) aligned / total accuracy (%) GPT-4-turbo 1418 / 2071 68.47 1438 / 2070 69.47 286 / 962 29.73 gemini-2.0-thinking 1354 / 2070 65.41 1621 / 2070 78.31 127 / 962 13.20 o1-mini 1444 / 2070 69.76 1401 / 2071 67.65 227 / 963 23.57 o3-mini 1448 / 2004 72.26 1206 / 2004 60.18 399 / 943 42.31 deepseek r1 1369 / 2071 66.10 1342 / 2071 64.80 173 / 964 17.95
Table 3. The evaluation results of each human label in LLMEval2. Only valid responses are counted when calculating the accuracy, while samples that couldn’t receive responses due to triggering Azure OpenAI’s content management policy are excluded. So there are some differences in the total values of different models.
@@ 5.2.3. Evaluation of Reasoning LLM-as-a-Judge
As discussed in Sections 2.4 and 2.5, judgment serves as the foundation for effective reasoning capabilities. In other words, models with stronger reasoning capabilities are generally better equipped to perform as reliable judges. To validate this assumption, we conducted evaluations on several reasoning LLMs including o1-mini, o3-mini, Gemini-thinking and Deepseek-R1. The results in Tables 2 and 3 provide key insights into the performance of reasoning-focused LLMs. While these models—gemini-2.0-thinking, o1-mini, o3-mini and deepseek r1—demonstrate competitive alignment and accuracy relative to the top-performing GPT-4-turbo, their improvements in tasks requiring human alignment are not as pronounced as expected.
GPT-4-turbo remains the benchmark for alignment, achieving the highest accuracy rates of 68.47 Among reasoning-enhanced models, gemini-2.0-thinking shows strong performance in the human=model2 scenario, achieving an accuracy of 78.27 These results indicate that reasoning-enhanced LLMs provide meaningful advancements over baseline models but fall short of delivering consistent advantages in alignment-related tasks, suggesting room for further optimization in this area.
@@ 6. LLM-as-a-Judge and o1-like Reasoning Enhancement
When faced with a challenging problem, humans often spend considerable time and effort reflecting on various possibilities before arriving at a solution. In a similar fashion, o1, an advanced model developed by OpenAI, engages in a structured chain of thought to solve complex tasks. This process of deliberate reasoning allows o1 to continuously refine its approach (Qin et al., 2024b), step by step, as it navigates through difficult scenarios. A key factor in enhancing o1’s reasoning is the integration of LLM-as-a-Judge, which evaluates the model’s reasoning paths at each stage. As o1 works through a problem, the judge provides feedback that helps the model improve by pointing out inconsistencies, suggesting corrections, and identifying simpler ways to break down difficult tasks. By using the feedback from its own evaluations, much like a constitutional AI framework, o1 is able to adapt its reasoning strategies and enhance its performance. Through reinforcement learning, o1 fine-tunes its strategies, learning not only from its successes but also from its mistakes. The combination of LLM-as-a-Judge, reinforcement learning, and the feedback loop from constitutional evaluation allows o1 to dynamically adjust its reasoning, ensuring that the model continuously improves its ability to solve complex problems over time. This synergy between reasoning and judgment, combined with continuous feedback, drives o1’s advanced problem-solving capabilities.
@@
Reasoning, as a cognitive process, involves applying logic and evidence to derive conclusions. It is central to intellectual tasks such as decision-making, problem-solving, and critical analysis. Reasoning requires evaluating multiple possibilities and determining the most logically sound and coherent path. In contrast, LLM-as-a-Judge refers to the use of LLMs to perform judgment tasks, such as evaluating, scoring, ranking, or selecting optimal answers based on generated outputs. This concept parallels a judge’s role in ensuring fairness, accuracy, and coherence in competitive contexts. Although reasoning and judgment are separate concepts, they are closely connected. As shown in Figure 15, reasoning frequently relies on judgment to evaluate intermediate steps, improve logic, and guarantee clarity in its results. When the process involves an infinite number of judgments, we can consider it as a process that approximates both reasoning and thinking. At the same time, effective judgment relies on strong reasoning capabilities to evaluate options against a set of logical criteria, which we discussed in Section 2.5 and Section 5.2.3. Therefore, LLM-as-a-Judge not only evaluates outputs but also enhances the reasoning process by helping to identify the most coherent, accurate solutions.
@@ 7. Applications
LLMs’ abilities as evaluators have gained widespread recognition in specialized fields, especially in complex, qualitative areas like legal texts, mathematical reasoning, and scientific research (Zhao et al., [n. d.]). This section reviews recent developments in LLM-as-a-judge applications across finance, law, science, and other industries, investigating how domain knowledge and LLM evaluators can further expand their impact in critical areas.
@@ 7.1.1. NLP
Text Generation Text generation tasks, such as dialog response generation, summarization, story creation, and creative writing, require content that is safe, accurate, and contextually relevant, though there isn’t a single ”correct” answer (Badshah and Sajjad, 2024; Bermejo, 2024). Unlike traditional metrics-based evaluations, LLM-as-a-judge offers a nuanced, adaptable, and customized assessment. According to Zheng et al. (2023b), LLMs like GPT-4 can evaluate text generation comparably to humans. This method has been used to evaluate outputs from single models and to compare multiple models in competitive settings. For instance, Gao et al. (2023b) employ ChatGPT for human-like summarization evaluation, while Wu et al. (2023) propose a comparison-based framework where LLMs act as judges to evaluate summarization quality.
@@
Reasoning Enhancing the reasoning capabilities of LLMs can overcome the limitations of scaling laws, unlocking their full potential. Effective reasoning is essential for tackling complex problems, making informed decisions, and delivering accurate, context-aware responses. Wei et al. (2022) introduce Chain-of-Thought (CoT) prompting to facilitate step-by-step reasoning. More sophisticated cognitive structures (Yao et al., 2023a; Hao et al., 2023) have been proposed to further enhance reasoning, yet selecting a reliable reasoning path remains a significant challenge. LLM-as-a-judge has been employed to address this issue.
@@
Some studies focus on sample-level reasoning path selection. Gao et al. (2023a) present a strategy evaluator for assessing candidate strategies. Kawabata and Sugawara (2024) propose REPS (Rationale Enhancement through Pairwise Selection), which uses pairwise self-evaluation to select valid rationales. Lahoti et al. (2023) demonstrate that LLMs can identify and enhance response diversity by aggregating multiple critiques. In multi-agent frameworks, Liang et al. (2023) introduce multi-agent debating (MAD), where a judge LLM selects the most reasonable response. Similarly, Li et al. (2024c) utilize a judge LLM in layer-based multi-agent collaboration to improve response quality and efficiency.
@@
For step-level reasoning path selection, LLMs act as process reward models (PRMs) to evaluate state scores. Creswell et al. (2023) break down reasoning into Selection and Inference, using LLMs to judge potential reasoning traces. Xie et al. (2024b) propose the Kwai-STaR framework, transforming LLMs into state-transition reasoners for mathematical reasoning. Lightman et al. (2023) train LLMs as PRMs for inference-time supervision and best-of-N sampling. Setlur et al. (2024) introduce process advantage verifiers (PAVs) to generate rewards based on the likelihood of future correct responses. Advanced cognitive structures are also simulated; Hao et al. (2023) use LLMs as a world model with Monte Carlo Tree Search (MCTS) for deliberate path selection. Besta et al. (2024) model LLM outputs as graphs to evaluate coherence and logical reasoning. Additionally, critique-based LLM judges (Ankner et al., 2024; Lan et al., [n. d.]; Yao et al., 2024; Yuksel and Sawaf, 2024) provide detailed feedback to enhance the reasoning process.
@@
Yao et al. (2023b) pioneered the use of LLMs in an interleaved manner to generate reasoning traces and task-specific actions. Reasoning traces guide the model in updating action plans, while actions facilitate interaction with external sources. Building on this, Yang et al. (2023) introduced Auto-GPT, which leverages LLM-as-a-judge to enhance tool usage accuracy. By integrating a variety of external tools, LLMs become more versatile, improving planning performance through judicious tool selection. Sha et al. (2023) explored the potential of LLMs in decision-making for complex autonomous driving scenarios, requiring human-like commonsense reasoning. Zhou et al. (2024b) employed a self-discovery process where LLMs judge queries and select the most suitable reasoning structure for subsequent inference.
@@
Retrieval The role of LLM-as-a-judge in retrieval encompasses both traditional document ranking and dynamic Retrieval-Augmented Generation (RAG) approaches. In traditional retrieval, LLMs enhance ranking accuracy through advanced prompting techniques, enabling effective document ordering with minimal labeled data. RAG frameworks leverage LLMs’ ability to generate content guided by retrieved information, supporting applications requiring complex or evolving knowledge integration.
@@
Recent studies have explored LLMs as judges for document ranking, aiming to boost precision and reduce reliance on extensive training data. Zhuang et al. (2024a) embed fine-grained relevance labels within LLM prompts, enabling models to distinguish subtle relevance variations for refined document ordering. Innovations in listwise ranking include Ma et al. (2023)’s Listwise Reranker with a Large Language Model (LRL), which reorders document identifiers without task-specific training data. Zhuang et al. (2024b) introduce a Setwise prompting strategy for zero-shot ranking, enhancing efficiency without sacrificing performance. To address positional biases, Tang et al. (2024c) propose permutation self-consistency, averaging multiple list orders to yield order-independent rankings. Qin et al. (2024a) critique pointwise and listwise ranking prompts, proposing Pairwise Ranking Prompting (PRP) with medium-sized, open-source LLMs as a cost-efficient alternative to larger models.
@@
Domain-specific RAG systems reveal LLMs’ potential to navigate complex queries by integrating specialized knowledge structures. Wang et al. (2024d) present BIORAG, enhancing vector retrieval with hierarchical knowledge structures and a self-aware evaluated retriever. Li et al. (2024d) introduce DALK, combining an LLM with a continuously evolving Alzheimer’s Disease knowledge graph, using self-aware knowledge retrieval for noise filtering. Jeong et al. (2024) propose Self-BioRAG, adapting RAG principles to biomedical applications Liu et al. (2023a), with LLMs selecting the best evidence for answer generation.